# MultiGA: Leveraging Multi-Source Seeding in Genetic Algorithms

### v1

@@ 4.3 Text-to-SQL ($\mathcal{D}_{\texttt{SQL}}$)
Text-to-SQL is a critical task for enabling non-technical users to query databases and extract insights without writing code. In industry, developing cost-efficient and accurate text-to-SQL frameworks is essential for building robust agentic systems. For this experiment, we used the BIRD mini-dev dataset, which contains 500 realistic industry-level questions paired with gold-standard SQL queries (li_can_2023). We partitioned the data into training and test sets, using the training set to construct positive and negative examples for each test query. Specifically, we embedded the training data and performed cosine similarity search to retrieve relevant natural language–SQL pairs for the current question. All prompting logic and retrieval configurations were defined in a task-specific configuration file to ensure reproducibility (see Appendix A).
@@ 4.4 Meeting Planning ($\mathcal{D}_{\texttt{NATPLAN}}$)
We next applied MultiGA to structured reasoning through the meeting scheduling benchmark introduced by zheng_natural_2024. In this task, the LLM must generate a plan that maximizes the number of valid meetings during a hypothetical trip to San Francisco. Solving it requires navigating multiple logistical constraints, such as travel distance between meeting locations, meeting durations, and participant availability. This benchmark thus provides a natural setting to test whether iterative recombination can improve complex structured outputs (see Appendix B).
4.5 GPQA Science Questions ($\mathcal{D}_{\texttt{GPQA}}$)
To broaden the evaluation of MultiGA, we tested the framework on the GPQA benchmark of graduate-level science multiple-choice questions (rein_gpqa_2023). This dataset covers advanced scientific domains and provides a challenging benchmark beyond code generation and optimization tasks. We used the Diamond subset, which contains the highest-quality questions in the benchmark (see Appendix C).
@@ 4.6 BBQ Bias Evaluation ($\mathcal{D}_{\texttt{BBQ}}$)
Finally, we evaluated whether MultiGA mitigates social bias using the BBQ benchmark (parrish_bbq_2022). We focused on the race and ethnicity subset, which contains over 6,000 examples. Each example presents an ambiguous or unambiguous scenario involving racial stereotypes, followed by a question designed to reveal whether the model exhibits biased behavior. Rather than using the benchmark in its standard multiple-choice format, we adapted it to an open-ended setting and subsequently extracted answers for accuracy evaluation (see Appendix D).
5 Results and Discussion
We evaluate accuracy on four benchmarks: Text-to-SQL ($\mathcal{D}_{\texttt{SQL}}$), GPQA Science Questions ($\mathcal{D}_{\texttt{GPQA}}$), Meeting Planning ($\mathcal{D}_{\texttt{NATPLAN}}$), and BBQ Bias ($\mathcal{D}_{\texttt{BBQ}}$ ). For each task, we compare seeding MultiGA with outputs from individual LLMs against seeding with all five models simultaneously ($\mathcal{G}$). Results are shown in Table 2.
Seed Model $\mathcal{D}_{\texttt{SQL}}$ $\mathcal{D}_{\texttt{NATPLAN}}$ $\mathcal{D}_{\texttt{GPQA}}$ $\mathcal{D}_{\texttt{BBQ}}$
gpt-4o 0.56 0.39 0.44 1.00
qwen2.5-coder-32b-instruct 0.45 0.28 0.30 0.87
deepseek-r1-distill-qwen-32b 0.46 0.23 0.41 0.97
openai/gpt-oss-20b 0.48 0.19 0.41 0.93
mistral-7b-instruct-v0.3 0.25 0.10 0.22 0.87
$\mathcal{G}$ (MultiGA) 0.55 0.40 0.40 1.00
Table 2: Accuracy across tasks when seeding with each individual model versus seeding with all five simultaneously ($\mathcal{G}$).
Across all four benchmarks, we observe that seeding with all models simultaneously ($\mathcal{G}$ ) produces accuracy that consistently approaches the performance of the strongest single-model baseline.
5.1 Meeting Planning ($\mathcal{D}_{\texttt{NATPLAN}}$)
For $\mathcal{D}_{\texttt{NATPLAN}}$, the $\mathcal{G}$ configuration not only avoids collapse from weaker models but slightly outperforms the top single seed by about 1%. We can see that this task requires multi-step temporal and spatial reasoning, integrating participant constraints, availability, and task ordering. Because prior state-of -the-art work has prioritized Gemini models for this benchmark, we also tested the MultiGA framework seeded only with gemma3-27b. However, this resulted in 29% accuracy compared to $\mathcal{G}$ with 40%. This result demonstrates MultiGA’s advantage in ambiguous settings where multiple solutions may be valid. This aligns with the broader motivation for genetic algorithms as optimization frameworks since MultiGA can iterate over potential solutions, prune out worse choices, and converge to a more optimal result.
5.2 Text-to-SQL ($\mathcal{D}_{\texttt{SQL}}$) and GPQA ($\mathcal{D}_{\texttt{GPQA}}$)
For $\mathcal{D}_{\texttt{SQL}}$, the $\mathcal{G}$ configuration (0.55) nearly matches the best single seed, gpt-4o (0.56). This task requires precise grounding in database schemas, where even small errors in column names or join conditions lead to incorrect queries. MultiGA narrows the gap to gpt-4o while providing stability against weaker models like mistral-7b (0.25) that often hallucinate schema elements. On GPQA, $\mathcal{G}$ (0.4015) falls within 0.03 of the best single seed, gpt-4o (0.4369). Since the task involves refining multiple-choice answers, there is less room for MultiGA to improve beyond single-LLM predictions. As expected, the gains here are more modest compared to the other benchmarks.
5.3 BBQ Bias Evaluation ($\mathcal{D}_{\texttt{BBQ}}$)
Finally, $\mathcal{D}_{\texttt{BBQ}}$ evaluates social bias, where weaker models often amplify stereotypes. Here, $\mathcal{G}$ achieves perfect accuracy (1.00), matching gpt-4o. Importantly, MultiGA demonstrates that combining weaker seeds does not degrade robustness on fairness-sensitive evaluation. For example, while Mistral-7b reached only 87% accuracy by itself, $\mathcal{G}$ pruned such biased outputs and instead preserved the judgments of less biased models through its selection mechanism. These results underscore the importance of diverse population initialization for tasks sensitive to demographic fairness, with gpt-4o-mini acting as an effective neutral evaluator and recombination agent, rather than a corrector of suboptimal outputs from weaker LLMs.
5.4 Robustness and Consistency
Compared to approaches such as Mind Evolution, MultiGA demonstrates a distinct strength: consistency. Mind Evolution leverages genetic algorithms to exploit inference-time compute, enhancing LLM reasoning through random exploration and convergent thinking (lee_evolving_2025). Its reported results are impressive, with near-perfect accuracy on tasks such as Trip Planning, Meeting Planning, and TravelPlanner when powered by Gemini 1.5 Pro and Flash (zheng_natural_2024; xie_travelplanner_2024). However, when evaluated under the same setup with gpt-4o-mini instead of Gemini, performance drops sharply—over 50% on Trip Planning, 20% on TravelPlanner, and 13% on Meeting Planning. Because reproducible code has not been released, it remains unclear whether these gaps generalize across models.
Our objective, by contrast, was not to surpass the state of the art on each benchmark, but to examine whether combining multiple LLMs could mitigate brittleness and large performance fluctuations. With a lightweight and transparent implementation, MultiGA generalizes across domains without relying on the strengths of a single model. Notably, when we tested replacing Mistral-7B-Instruct with GPT-5 for $\mathcal{G}$ on the meeting planning task, accuracy decreased from 40% to 36%. This result highlights that Mistral’s value lay in its diversity rather than raw performance, even though it was the weakest model individually. More broadly, it demonstrates that exchanging a smaller model with a larger, more expensive model is not required for MultiGA to achieve strong and reliable results, underscoring its efficiency as well as its effectiveness.
Furthermore, MultiGA is not expected to consistently outperform the best individual seed, since the initial population may include far weaker candidates that must be discarded. Yet its consistent ability to converge toward near-best performance highlights how evaluator-guided selection and recombination amplify strong solutions while eliminating poor ones. This makes MultiGA a robust choice when the most suitable model for a task is unknown or less accessible, adaptively closing the gap to the strongest baseline without manual tuning.
Limitations
While MultiGA demonstrates the potential of combining the perspectives of different models, some limitations remain that suggest directions for future work. First, our evaluation focuses on a limited set of benchmarks—Text-to-SQL, Meeting Planning, GPQA Science, and BBQ Bias—which, although diverse, do not capture the full range of reasoning and generative tasks. Expanding to additional evaluations such as SuperGLUE or multi-hop reasoning datasets could further test the framework’s generality (wang_superglue_2020; yang_hotpotqa_2018). We also wish to increase dataset sizes and explore how deeper generational runs correlate with population diversity and task complexity.
Second, we primarily used gpt-4o-mini as the evaluator for consistency and cost efficiency. Although effective for evaluation and recombination, different evaluator architectures or mixtures of evaluators could lead to different evolutionary trajectories. Evaluator choice therefore remains a potential source of bias, and exploring alternative evaluators may improve our understanding of multi-source genetic optimization. Despite these constraints, MultiGA’s transparency and modularity make it a promising framework for orchestrating heterogeneous LLMs toward more reliable reasoning.
@@ 6 Conclusion
Our experiments across four distinct benchmarks—text-to-SQL generation, complex meeting planning, graduate-level scientific reasoning, and social bias evaluation demonstrate that MultiGA is a robust and effective strategy. The framework consistently converges to a solution that rivals the accuracy of the best-performing individual model. By systematically pruning weaker solutions and semantically recombining the strengths of stronger ones, MultiGA provides a reliable method for achieving high-quality results without requiring prior knowledge of which model is best suited for a particular task.
Future studies could investigate the impact of the evaluator model’s capabilities on the evolutionary process or explore dynamic seeding strategies where the pool of generator LLMs is adapted over time. Finally, extending the MultiGA framework to more complex, multi-step agentic workflows could unlock new possibilities, allowing different models to contribute specialized strengths at various stages of a problem-solving process.
Appendix A SQL Task: System Instructions and Prompts
A.1 Prompts for Generating Initial SQL Solutions
@@
init_sol_system = """You are a specialized SQL query generator. You receive pre-processed inputs from an upstream schema linking agent
and focus on generating accurate and executable SQL representations of a users natural language question."""
@@
Task: Generate a SQL Query based on the users query.
@@
Database Context: {db_schema}
@@
Your Focus: SQL Generation Excellence
Linking Question to DB
Use the linked schema elements as your primary source for matching the users question with the database schema
However, you must consult the actual database schema when semantic score is low
Query Construction Rules
ALWAYS use original_column_name NOT column_name from the database
Do not define aliases using the AS clause. Only use real names.
Try not to make queries that return irrelevant excess information
Edge Case Handling
If schema linking appears incomplete:
Use database context to fill gaps
Dont hallucinate or guess about column names
Pattern Learning
@@
Output Structure:
Return the SQL query only, no markdown formatting and DO NOT wrap in ‘‘‘sql
Example ouput (use this to see the required structure): {output_example}
DO NOT WRAP WITH TRIPLE BACKTICKS
A.2 Prompts for Crossover Phase
@@
crossover_system = """You are tasked to perform the crossover in a genetic algorithm aimed at creating a correct SQL query to match a users question. """
@@
crossover_prompt_unfilled = """ You are given two SQL queries, each attempting to express a users natural language question against a specific database.
Your task is to analyze both queries and synthesize a new SQL query that combines the strengths of each, mimicking the crossover operation in genetic algorithms.
The objective is to produce a child query that more effectively and accurately translates the users intent: {query}.
Here are the parent SQL queries:
@@
Database Context: {db_schema}
Supporting context for query construction
Reference for relationships and constraints
Domain Evidence: {evidence}
Business context for accurate query logic
Helps with aggregation and filtering decisions
Consider the following when generating the child SQL query:
Evaluate whether certain filters are necessary and relevant to the users intent
Ensure that if data is required from multiple tables, appropriate JOIN operations are included in the SQL query
Output Structure:
Return the SQL query only, no markdown formatting and DO NOT wrap in ‘‘‘sql
DO NOT WRAP WITH TRIPLE BACKTICKS """
A.3 Prompts for Objective Function (Evaluation Phase)
Listing 5: System and feedback prompts for SQL objective function evaluation.
system_instructions = """Given a users question and a SQL query, you are tasked to determine how well the SQL captures
the meaning of the question and appropriately curates the SQL query with respect the given database. """
@@ A.3 Prompts for Objective Function (Evaluation Phase)
Task: Evaluate the correctness of the following SQL query in capturing the intent of the users question {user_query}.
You are given a SQL Query below and the output (which may be truncated) of querying it on the appropriate database:
@@
Output: {output}
More helpful context and information is listed below:
Database Context: {db_schema}
Supporting context for query construction
Reference for relationships and constraints
Domain Evidence: {evidence}
Business context for accurate query logic
Helps with aggregation and filtering decisions
Reference Date: {current_date}
Your Focus: Quality Assessment
Compare the users question, the generated SQL query, and the outcome of the SQL query on the database.
If there is an error in the SQL output, that is an indication of a poorly crafted query.
If there is not an error, ensure that the evidence is being correctly applied, and that all filters and conditions are actually needed.
Also please check the source table for each column and ensure proper JOIN operations are included wherever required.
You may find the following SQL rules helpful.
GROUP BY or ORDER BY statements
If a variable is used in the GROUP BY or ORDER BY clause, it must also appear in the SELECT statement.
You should only group or order by a column if it is included in the SELECT statement.
JOINS
Ensure that if you are taking information from MORE than one table, you are doing a JOIN! This is a common mistake where LLMs are forgetting to do JOINs.
Always specify the join condition explicitly using ON table1.column = table2.column and avoid Cartesian products.
Rate the SQL query on a scale from 0.00 to 1.00, where 1.00 represents a perfect match to the natural language question and 0.00 represents a very poor match.
Example Outputs:
A nearly perfect SQL translation with only a minor flaw -> 0.90
An incorrect SQL translation that barely matches the question -> 0.15
DO NOT INCLUDE OTHER TEXT. Please only return the float in your output.
Appendix B Meeting Planning Task: System Instructions and Prompts
B.1 Prompts for Generating Initial Solutions

### v2

@@ 5.1 Text-to-SQL ($\mathcal{D}_{\texttt{SQL}}$)
Text-to-SQL tasks require models to closely follow the structure of the underlying database schema, so refinement or corrective methods can be valuable. The jump in accuracy from the 0-shot and 3-shot prompting approaches to the MultiGA framework demonstrates the power of this refinement method for complex multi-step tasks (See Table 2). MultiGA effectively combats schema alignment issues and provides stability against weaker models like gemma-2 (0.475) by pruning out incorrect text-SQL translations. On $\mathcal{D}_{\texttt{SQL}}$, our $\mathcal{G}_{\text{A}}$ configuration achieves 0.705 accuracy on 200 test rows, substantially outperforming the reported GPT-4 baseline of 0.478 on the full 500-row set and exceeds the TA + GPT-4o approach (0.630) by over 7 percentage points. [^3] Reported scores from the BIRD Mini-Dev benchmark: https://github.com/bird-bench/mini_dev Additionally, we can see that when seeding the framework with GPT-5 only, the accuracy drops by 12 percent, highlighting that a single large model may perform worse on a multi-stage task compared to a collection of potentially smaller models than compliment one another. Even when seeded with Llama-4 alone, the MultiGA framework outperforms seeding with only GPT-5, and also beats the GPT-4 baseline mentioned previously.
@@ 5.2 Meeting Planning ($\mathcal{D}_{\texttt{NATPLAN}}$)
For $\mathcal{D}_{\texttt{NATPLAN}}$, the configuration using GPT-5 as the sole population seed and Qwen3 as the evaluator achieved a the highest score (0.95) under the benchmark’s accuracy metric. Importantly, this metric evaluates only whether the generated output matches the number of valid meeting plans, rather than the exact plan itself. [^4] Meeting Planning Evaluation: https://github.com/google-deepmind/natural-plan $\mathcal{G}_{\text{A}}$ (0.88) did not attain the highest performance across the experiments using our genetic algorithm, though it still outperforms the 0-shot and 5-shot baselines. This suggests that when a single model already excels at a task, the benefit of multi-source seeding diminishes, as weaker candidate solutions must first be pruned. By contrast, results on $\mathcal{D}_{\texttt{SQL}}$ indicate that increased task and pipeline complexity may favor greater model diversity. Additionally, there is a significant gap between $\mathcal{G}_{\text{A}}$ (0.880) and $\mathcal{G}_{\text{B}}$ (0.690). This performance gap may arise from differences in evaluator choice or model composition, and it shows how relatively small changes in model selection could lead to large accuracy differences on certain tasks. In this meeting planning task, there is much freedom for the LLM to explore different planning options, resulting in varied responses and accuracy.
5.3 GPQA ($\mathcal{D}_{\texttt{GPQA}}$)
On the Diamond subset of the graduate level science question benchmark, $\mathcal{D}_{\texttt{GPQA}}$, the $\mathcal{G}_{\text{A}}$ configuration attains an accuracy of 0.959, which was the highest accuracy seen across our experimental configurations, and also exceeds the Gemini 3 Pro baseline (0.917). [^5] GPQA benchmark leaderboard: https://www.vals.ai/benchmarks/gpqa $\mathcal{G}_{\text{B}}$ also surpasses this baseline at 0.919, indicating that both evaluator choices yield strong performance. One possible explanation is that in multiple-choice settings, the discrete solution space limits the impact of $E$, since recombination largely reduces to selection among the provided answer options. It is worth to note that $\mathcal{G}_{\text{A}}$ relies primarily on open-source models, as Qwen3 as $E$ does all of the recombination and evaluation work in the pipeline, yet it achieves the highest accuracy, demonstrating that competitive performance on QA benchmarks can be achieved with more accessible models when utilizing the MultiGA framework.
@@ 5.4 BBQ Bias Evaluation ($\mathcal{D}_{\texttt{BBQ}}$)
[…]
@@ 6 Conclusion
We run experiments across four distinct benchmarks—text-to-SQL generation, meeting planning, graduate-level scientific reasoning, and social bias evaluation. This inference-time optimization framework produces strong results and even surpasses existing public benchmarks in certain tasks ($\mathcal{D}_{\texttt{SQL}}$, $\mathcal{D}_{\texttt{GPQA}}$).
MultiGA demonstrates two key strengths: strong performance and robustness across model families. Our objective is not to surpass the state of the art on each task, but to examine whether combining multiple LLMs can mitigate brittleness and large performance fluctuations. In particular, MultiGA shows that using a larger, more expensive model throughout an entire experimental pipeline is not required to achieve strong and reliable results, highlighting both its efficiency and its ability to extract high-level insights from more capable models before relying on more accessible ones.
Moreover, MultiGA is not expected to consistently outperform the strongest individual seed, as the initial population may include substantially weaker candidates that must be filtered out. Nevertheless, its convergence toward near-best or top performance highlights its effectiveness in amplifying high-quality solutions while eliminating poor ones. This systematic pruning provides a reliable method for achieving high-quality results without prior knowledge of which model is best suited for a particular task.
Future studies could more thoroughly investigate the impact of the evaluator model’s capabilities on the evolutionary process or explore dynamic seeding strategies where the pool of generator LLMs is adapted over time. Finally, extending the MultiGA framework to more complex, multi-step agentic workflows could unlock new possibilities, allowing different models to contribute specialized strengths at different stages of a problem-solving process.
Appendix A Appendix
Figure 3: Example of recombination on a text-to-SQL task (“Among the patients who have an abnormal level of glutamic oxaloacetic transaminase, when was the youngest of them born?”). Parent 1 (Llama-4-Maverick-17b) and Parent 2 (Gemma-2-27b) are recombined by Qwen3 to produce a child solution that achieves a perfect accuracy score of 1.0, matching the ground-truth result.
Appendix B SQL Task: System Instructions and Prompts
B.1 Prompts for Generating Initial SQL Solutions
@@
init_sol_system = """You are a specialized SQL query generator for SQLite databases. You receive pre-processed inputs from an upstream schema linking agent and focus on generating accurate and executable SQLite-compatible SQL representations of a users natural language question.
Your goal is to provide clear, step-by-step reasoning that demonstrates your understanding of the database schema and query requirements, followed by your final SQL query.
CRITICAL: You must use SQLite syntax only. SQLite does not support functions like YEAR(), MONTH(), DAY(). Instead use strftime()."""
@@
Task: Generate a SQLite-compatible SQL Query based on the users query.
IMPORTANT: This is a SQLite database. Use SQLite syntax only:
- Date/time functions: Use strftime(’%Y’, date) NOT YEAR(date)
- Date extraction: strftime(’%Y’, date) for year, strftime(’%m’, date) for month, strftime(’%d’, date) for day
- No CURDATE(), NOW() - use date(’now’), datetime(’now’)
@@
Database Context (SQLite): {db_schema}
@@
# Instructions
Think through the problem systematically before writing SQL:
1. Identify what information the user is asking for
2. Map the question to relevant tables and columns
3. Determine necessary JOINs based on table relationships
4. Apply appropriate filters, aggregations, and ordering
5. Verify SQLite compatibility of all functions used
# Pattern Learning
@@ 5. Verify SQLite compatibility of all functions used
Provide your solution in this format:
## Reasoning
[Your step-by-step analysis]
- Key entities identified from the question
- Tables and columns needed
- JOIN strategy (if multiple tables)
- Filters and conditions required
- Aggregation/grouping logic (if needed)
- SQLite-specific considerations
## SQL Query
[Your final SQL query - no markdown formatting, no triple backticks]
B.2 Prompts for Crossover Phase
@@
crossover_system = """You are a creative and skilled engineer who knows everything about SQLite database designs. You are tasked to perform the crossover in a genetic algorithm aimed at creating a correct SQLite-compatible SQL query to match a users question.
Your goal is to analyze the reasoning and SQL from both parent solutions, identify their strengths and weaknesses, and synthesize an improved solution with clear reasoning.
A crossover combines the best reasoning and SQL patterns from both parents while avoiding their mistakes. Remove unnecessary complexity and focus on correctness.
CRITICAL: Only use SQLite syntax. Never use YEAR(), MONTH(), DAY() functions - use strftime() instead."""
@@
crossover_prompt_unfilled = """You are performing a crossover operation in a genetic algorithm for SQL query generation.
# Users Question:
{query}
# Parent Solution 1:
@@
# Database Context (SQLite):
{db_schema}
# Domain Evidence:
{evidence}
# Feedback from Previous Attempts:
{memory_context_section}
# Instructions
1. Review the reasoning chains from both parents carefully
2. Identify the strongest SQL patterns and logic from each
3. Combine valid reasoning steps - take the best from both
4. If parents disagree on approach, use database schema and evidence to resolve
5. Learn from feedback: avoid patterns that scored poorly
6. Build on patterns that scored well
7. CRITICAL: Use SQLite functions only (strftime, NOT YEAR/MONTH/DAY)
IMPORTANT: If previous generations show errors like "no such function: YEAR", avoid that mistake!
Provide your synthesized solution:
## Reasoning
[Combined reasoning drawing from both parents strengths]
- Key entities and their mapping to tables
- Analysis of parent approaches - what worked, what didnt
- Resolution of any contradictions between parents
- JOIN strategy and filter logic
- Final justification for your approach
## SQL Query
[Your final SQL query - no markdown formatting, no triple backticks]
B.3 Prompts for Objective Function (Evaluation Phase)
Listing 5: System and user prompts for SQL objective function evaluation.
system_instructions = """You are an expert evaluator assessing the quality of SQL query reasoning.
IMPORTANT: You are evaluating REASONING QUALITY ONLY, not whether the query produces correct results.
Do NOT try to determine if the output is correct - focus purely on the reasoning process and SQL construction quality."""
@@ B.3 Prompts for Objective Function (Evaluation Phase)
Task: Evaluate the REASONING QUALITY of the following SQL solution.
# Users Question:
{user_query}
# Database Context (SQLite):
{db_schema}
# Domain Evidence:
{evidence}
# Candidate Solution:
{candidate_response}
# SQL Query Extracted:
@@
# Execution Result:
{output}
# Evaluation Criteria - REASONING QUALITY ONLY
Evaluate the SQL reasoning process (ignore whether the output is correct):
- **5**: Excellent - Clear logical progression, correct schema understanding, appropriate JOINs and filters
- **4**: Good - Sound reasoning with minor gaps, demonstrates solid SQL understanding
- **3**: Adequate - Generally correct approach but with some logical issues or missing steps
- **2**: Weak - Shows some understanding but significant reasoning flaws or gaps
- **1**: Poor - Minimal valid reasoning, mostly incorrect logic or very shallow analysis
- **0**: Invalid - No coherent reasoning or completely wrong SQL approach
Consider:
- Quality of schema analysis and table selection
- Logical coherence of JOIN strategy
- Appropriate use of filters and conditions
- SQLite compatibility (strftime vs YEAR/MONTH/DAY)
- Clarity of reasoning steps
- Depth of analysis (not just surface-level)
# Output Format (follow exactly):
REASONING_SCORE: [0-5]
REASONING_FEEDBACK:
[What was good/bad about the SQL reasoning? Be specific about schema understanding and query logic.]
IMPROVEMENT_SUGGESTIONS:
[How could the reasoning be strengthened? What SQL patterns or concepts are missing?]
Appendix C Meeting Planning Task: System Instructions and Prompts
C.1 Prompts for Generating Initial Solutions